i am having an extremly hard time fuguring this out can someone please help
solve by the substitution method
5x-2y=-19
5x +27=y



Answer by ankor@dixie-net.com(22740)

5x - 2y =-19
5x + 27 = y
:
Here it is step-by-step:
:
You can see from the 2nd equation that y = (5x+27); therefore:
Replace y with (5x+27) in the 1st equation.
5x - 2(5x+27) = -19
5x - 10x - 54 = -19
5x - 10x = -19 + 54
Combine like terms
-5x = +35; corrected a mistake here
x has to be positive, multiply both sides by -1
5x = -35
x = -35/5
x = -7
:
We can find y using the 2nd equation y = 5x + 27
y = 5(-7) + 27
y = -35 + 27
y = -8
;
:
Confirm this by replace x and y in the 1st equation 5x-2y=-19:
5(-7) - 2(-8) = -19
-35 - (-16) = -19
-35 + 16 = -19; minus a minus is a plus
-19 = -19
:
Our solution: x=-7, y=-8